### v1

@@ 3.3 Quantum Architecture Search Methods
We summarize the various QAS algorithms in Table 1. The table organizes the characteristics of QAS algorithms, particularly focusing on the methods for updating structural parameters and whether the updates of structural parameters and rotation parameters are based on gradient descent. Additionally, we indicate whether super-circuits are required during the search process and whether noise experiments are included. The specific algorithm classifications are as follows:
Table 1: Review of quantum architecture search (quantum logic circuit synthesis) methods. “End-to-end” means the updating of both circuit structure and parameters are done by gradient descent. “Super-circuit” stands for pre-fixed circuit structure using human knowledge.
@@
Paper Method Application Sourced training Structure Updating Method Rotation Parameter Updating Super-circuit Free End-to-End Circuit Noise
@@
3.3.1 Heuristic Algorithms
Heuristic algorithms, particularly Genetic Algorithms (GAs), were among the earliest methods applied to automate quantum logic circuit synthesis [192]. Initial research focused on using GAs to search for simple circuits, such as quantum teleportation. Subsequent studies demonstrated the potential of GAs in evolving relatively simple quantum circuits, including quantum error correction codes and quantum adders [125, 15, 104, 103, 154]. The basic process of GAs involves: 1) selecting the initial population; 2) evaluating individuals in the current population; 3) sampling based on individual quality to form a candidate pool; and 4) performing crossover and mutation in the candidate pool. Steps 2 to 4 are repeated until a stop condition is met.
GAs are particularly suitable for problems where the solution is a sequence, as sequences can be viewed as strands of genes. From the quantum circuit representation methods introduced in the previous section, we can see that quantum circuits can be represented as sequences, with each gene locus corresponding to a quantum gate. Information such as gate type, parameters, and qubit index can be encoded as integer strings. The evaluation method in step 2 can refer to equations 9 and 10, while crossover and mutation correspond to modifying the position, type, parameters, and qubit information of quantum gates. The performance of GAs is influenced by factors such as initial population size, candidate pool size (number of sampled quantum circuits), and gene length (quantum circuit depth). Increasing the types of gates in the circuit further extends the encoding length for each quantum gate, adding complexity to the optimization process.
@@ 3.3.1 Heuristic Algorithms
3.3.2 Reinforcement Learning Algorithms
@@
To enhance the performance of QAS, researchers have explored machine learning algorithms beyond heuristic approaches, leveraging their powerful capabilities. Reinforcement learning (RL), as a representative example of unsupervised learning, has emerged as a suitable alternative, given the difficulty of introducing labeled input-output pairs in QAS. RL algorithms typically operate in discrete time steps, where an agent makes decisions at each step, and the environment provides corresponding rewards. The agent’s goal is to maximize rewards through a limited set of actions. In QAS, the agent’s action pool comprises candidate quantum gates, with each decision adding a gate to the circuit. The reward is calculated using a loss function after each addition. A key advantage of RL is its ability to generate circuits without preset maximum depths, theoretically allowing for continuous circuit deepening. Unlike heuristic methods, RL accommodates parameterized quantum gates, optimizing internal parameters while adding gates.
@@
Primitive studies [101, 202] tested RL algorithms on two-qubit Bell states and three-qubit GHZ states. RL algorithms are then applied to automate the synthesis of variational quantum circuits for solving ground state energies [147]. Their experiments with four-qubit and six-qubit lithium hydride Hamiltonians assigned a -1 reward for each added gate and a large positive reward for achieving chemical accuracy (i.e., $1.6\times 10^{-3}$ Ha). This reward mechanism successfully guided the agent to find circuits achieving chemical accuracy, validating RL’s feasibility. Recently, Patel et al. [150] introduced curriculum learning, enhancing the guidance mechanism towards chemical accuracy and demonstrating significant performance improvements over previous work.
3.3.3 Sampling-based Learning Algorithms
@@
In addition to reinforcement learning, sampling-based learning algorithms represent another significant class of circuit synthesis methods. These algorithms initially define the search space, considering a quantum circuit with $n$ qubits, $m$ layers, and $k$ candidate gates, resulting in a search space of size $k^{m\times n}$. The methodology proceeds as follows: first, sampling is performed on the $n\times m$ grid points to generate a complete quantum circuit. Subsequently, the sampled circuit undergoes evaluation, informing modifications to the sampling strategy for future iterations. This process is repeated iteratively to complete circuit synthesis. Notably, sampling-based learning algorithms demonstrate compatibility with parameterized quantum gates, allowing for optimization of internal circuit parameters post-sampling. The primary challenge inherent in this approach stems from the discrete nature of the sampling process, which precludes the production of differentiable gradients. Consequently, developing effective mechanisms for updating the sampling strategy remains a critical area of research in this domain.
Zhang et al. [208] proposed a neural network predictor for circuit synthesis, training it on small-scale quantum circuits and applying it to slightly larger-scale problems. This approach, while one of the few supervised learning models for QAS, faces limitations due to the diversity of quantum gates and the non-trivial relationship between gate configurations and circuit purposes. Recently, He et al. [71] introduced a training-free scheme based on [208], which reverted to sampling-based optimization. However, this study’s evaluation scale was limited (up to six qubits) and lacked comparison with state-of-the-art algorithms.
@@
Zhang et al. [209] employed Monte Carlo gradient estimation methods to update sampling strategies, demonstrating efficacy in Quantum Fourier Transform and error mitigation problems. Du et al. [48] selected the optimal circuit through multiple sampling and ranking, testing their algorithms on four-qubit hydrogen ground state energy prediction and image classification QNN design problems. Wang et al. [187] proposed a method similar to [48], focusing on noise adaptation for quantum hardware. This approach generates shallow quantum circuits under the constraints of the quantum hardware to mitigate errors. Experimental results on IBM quantum devices demonstrated significant improvements over manually designed deeper circuits. However, both [48] and [187] utilize pre-defined circuit modules (termed super-circuits or supernets), substantially reducing the search space through pre-designed modules or specified layer arrangements. This approach potentially limits the demonstration of the algorithms’ true capabilities, as successful circuit design relies heavily on pre-provided key circuit modules.
@@
Lu et al. [115] addressed this issue by questioning the efficacy of circuit structure sampling and synthesis and aligning algorithm applications. The study proposed using the arbitrary unitary approximation problem to benchmark QAS algorithms, considering this task the most challenging in QAS. To accommodate potentially deep decomposition depths, a simplified version - arbitrary circuit reconstruction - was introduced. This task provides a unitary matrix with a known circuit solution within a certain depth, enabling fairer algorithm comparisons. Results indicated that after removing prior information, such as circuit modules, the performance of [48] and [209] decreased significantly. Notably, when candidate gates were limited to non-parameterized quantum gates, machine learning algorithms generally underperformed compared to heuristic algorithms.
@@
To enhance the efficiency and accuracy of searching algorithms, Wu et al. [195] proposed a method utilizing Gumbel-Softmax [66, 20, 86] to bridge the gap between discrete sampling and continuous gradient updates. This method uses the argmax term in the forward pass and the softmax term in backpropagation, addressing issues inherent in sampling-based learning algorithms and significantly improving search efficiency. The study eliminates prior knowledge interference by searching directly within the complete $k^{m\times n}$ space. It evaluates the method on three problems: VQE circuits (ground state energy prediction and Max-Cut) and QNN circuits (image classification), demonstrating notable improvements. Furthermore, the article examines the scalability of QAS algorithms, proposing two search processes: Macro Search and Micro Search. Macro Search conducts global domain searches, while Micro Search uses preset sub-circuit modules with shared structural parameters but distinct internal rotation parameters. Unlike [48] and [187], which predetermine circuit module structures, Wu et al. [195] searches for the module structures. They facilitate the identification of frequently reused modules, such as single and double excitation operators in ground state energy estimation or quantum convolution and pooling modules in image classification. This approach reduces the search space and enables the application of modules discovered in small-scale problems (e.g., six-qubit lithium hydride) to larger-scale problems (e.g., 18-qubit methane). Given that [195] and [209] render the entire search process differentiable end-to-end, this approach can be categorized as a new track: differentiable quantum architecture search.
4 Quantum Logic Circuit Optimization
@@
In the context of quantum computing, several unavoidable factors must be considered when executing quantum circuits, including qubit coherence time, environmental noise, etc. These constraints necessitate the optimization of quantum logic circuits to reduce their complexity. While quantum logic circuit synthesis creates circuits from scratch, optimization algorithms refine existing circuits based on specific targets. Consequently, circuits produced by synthesis algorithms can be further enhanced through optimization techniques. Moreover, by integrating the constraints of circuit synthesis with the optimization targets, we can develop a unified synthesis optimization algorithm. This approach enables the direct creation of optimal quantum logic circuits that meet current optimization objectives.
4.1 Problem Definition
@@
We begin by presenting a formulation of the quantum logic circuit optimization problem. Analogous to the quantum logic circuit synthesis problem, we employ a unitary matrix to represent a given quantum circuit. Let $\mathbf{U}$ denote the unitary matrix of the original circuit to be optimized, and $\hat{\mathbf{U}}$ represent the unitary matrix of the optimized circuit. Given a set of optimization targets $\kappa_{i}$, where we assume that all targets are to be minimized, the corresponding optimization objectives can be expressed as:
$\min\sum_{i}\kappa_{i}(\hat{\mathbf{U}})$ (12)
subject $\text{subject to}\ ||\mathbf{U}-\hat{\mathbf{U}}||_{F}\leq\varepsilon.$
In this formulation, $\varepsilon$ denotes the permissible unitary matrix error. When $\varepsilon$ equals 00, the two circuits are required to be completely equivalent. The term $\kappa_{i}(\hat{\mathbf{U}})$ represents the score of the optimized circuit with respect to a specific optimization target. In the following sections, we will summarize and categorize the primary optimization objectives of quantum logic circuits, followed by a systematic review of various optimization methodologies.
4.2 Quantum Logic Circuit Optimization Targets
@@
The objectives of quantum logic circuit optimization algorithms are directly correlated to their optimization methods. Table 2 illustrates the optimization methods in various research and their corresponding optimization objectives. It can be seen that, excluding brute-force search and reinforcement learning, other methods are aimed at more specific optimization targets. Overall, there is an emphasis on optimizing gate count and CNOT gates, while optimization of T gates is closely related to phase, resulting in a comparatively limited array of available optimization methods.
@@
Table 2: Quantum logical circuit optimization methods versus optimization target.
@@ 4.2.1 Gate Count
Optimization of the number of gates represents a fundamental objective in quantum circuit design. This optimization remains crucial across the current NISQ era and prospective fault-tolerant quantum computing paradigms. In the NISQ context, where qubit coherence time is limited and noise effects are significant, minimizing the number of quantum gates serves to reduce overall circuit noise, thereby enhancing operational accuracy. Moreover, in the realm of fault-tolerant quantum computing, gate count reduction contributes to improved computational efficiency of quantum hardware. Thus, the imperative to minimize quantum gates persists as a key strategy for advancing quantum circuit performance, irrespective of the underlying quantum computing framework.
Various works such as [99, 199, 56, 110, 129, 14, 76, 122, 124, 123, 27, 166, 156, 126] focus on optimizing quantum gate count in different types of circuits. Methods proposed by [99, 156, 166] guarantee optimal Clifford circuits, with [166] demonstrating scalability. We also have works targeting other gate sets, such as the parameterized gate set [199], Toffoli networks [122], Multi-Control Toffoli (MCT) circuits [129, 14], and NCV circuits [124, 123]. There are also works focusing on reducing the gate count for VQE circuits [126].
4.2.2 Circuit Depth
@@
Circuit depth is intrinsically associated with the total runtime of a quantum circuit, which is constrained by the coherence time of qubits. Consequently, optimizing circuit depth holds practical significance for scalable quantum computing in the NISQ era. While reducing gate count and minimizing circuit depth are often considered alike, many studies address both objectives concurrently [99, 56, 110]. However, these goals are not entirely identical. Depth optimization aims to enhance circuit parallelism, thereby reducing overall execution time and improving accuracy and efficiency. A circuit with an optimal gate count does not necessarily imply optimal depth. For instance, a quantum circuit could be optimized to contain either 7 gates executed serially or 10 gates executed in 5 parallel layers. The former achieves the gate count optimization goal (7 gates), while the latter satisfies the depth optimization objective (5 layers). Therefore, this study distinguishes between these two optimization targets.
@@
Circuit depth optimization is typically achieved by increasing the parallelism of quantum gates. Theoretically, a set of quantum gates without dependencies on each other can be executed in parallel. Dependencies in this context commonly refer to overlapping qubit operations. Thus, depth optimization can be understood as partitioning all gates in a circuit into the minimum number of mutually independent gate groups. Numerous studies, including [99, 56, 110, 153, 74, 73, 32, 98, 8] have focused on circuit depth optimization. Specifically, Kliuchnikov et al.[99] achieved optimal circuits through recursion, a method that can be used to construct optimal circuit datasets. The algorithm presented in [8] extends beyond depth optimization with the capability of various goals. Other works focused on reducing the circuit depth of variational quantum circuits, especially QAOA circuits [74, 73, 32] and VQE circuits [98].
4.2.3 CNOT Count

### v2

@@ III-C Quantum Architecture Search Methods
We summarize the various QAS algorithms in Table I. The table organizes the characteristics of QAS algorithms, particularly focusing on the methods for updating structural parameters and whether the updates of structural parameters and rotation parameters are based on gradient descent. Additionally, we indicate whether super-circuits are required during the search process and whether noise experiments are included. The specific algorithm classifications are as follows:
TABLE I: Review of quantum architecture search (quantum logic circuit synthesis) methods.
@@
Paper Method Application Sourced training Structure Updating Method Rotation Parameter Updating Super-circuit1 Free End-to-End2 Circuit Noise
@@
  • 1
    pre-fixed circuit structure using human knowledge.       2 the updating of both circuit structure and parameters are done by gradient descent.
III-C1 Heuristic Algorithms
Those particularly Genetic Algorithms (GAs), were among the earliest methods applied to automate quantum logic circuit synthesis [125]. Initial research focused on using GAs to search for simple circuits, e.g. quantum teleportation. Subsequent studies demonstrated the potential of GAs in evolving relatively simple quantum circuits, including quantum error correction codes and quantum adders [131, 132, 133, 102, 134]. The basic process of GAs involves: 1) selecting the initial population; 2) evaluating individuals in the current population; 3) sampling based on individual quality to form a candidate pool; and 4) performing crossover and mutation in the candidate pool. Steps 2 to 4 are repeated until a stop condition is met.
GAs are particularly suitable for problems where the solution is a sequence, as sequences can be viewed as strands of genes. From the quantum circuit representation methods introduced in the previous section, we can see that quantum circuits can be represented as sequences, with each gene locus corresponding to a quantum gate. Information e.g. gate type, parameters, and qubit index can be encoded as integer strings. The evaluation method in step 2 can refer to equations 5 and 6, while crossover and mutation correspond to modifying the position, type, parameters, and qubit information of quantum gates. The performance of GAs is influenced by factors e.g. initial population size, candidate pool size (number of sampled quantum circuits), and gene length (quantum circuit depth). Increasing the types of gates in the circuit further extends the encoding length for each quantum gate, adding complexity to the optimization process.
@@ III-C1 Heuristic Algorithms
III-C2 Reinforcement Learning (RL) Algorithms
@@
To enhance the performance of QAS, researchers have explored machine learning algorithms beyond heuristic approaches, leveraging their powerful capabilities. RL, as a representative example of unsupervised learning, has emerged as a suitable alternative, given the difficulty of introducing labeled input-output pairs in QAS. RL algorithms typically operate in discrete time steps, where an agent makes decisions at each step, and the environment provides corresponding rewards. The agent’s goal is to maximize rewards through a limited set of actions. In QAS, the agent’s action pool comprises candidate quantum gates, with each decision adding a gate to the circuit. The reward is calculated using a loss function after each addition. A key advantage of RL is its ability to generate circuits without preset maximum depths, theoretically allowing for continuous circuit deepening. Unlike heuristic methods, RL accommodates parameterized quantum gates, optimizing internal parameters while adding gates.
@@
Primitive studies [128, 127] tested RL on two-qubit Bell states and three-qubit GHZ states. RL algorithms are then applied to automate the synthesis of variational quantum circuits for solving ground state energies [109]. Their experiments with four-qubit and six-qubit lithium hydride Hamiltonians assigned a -1 reward for each added gate and a large positive reward for achieving chemical accuracy (i.e., $1.6\times 10^{-3}$ Ha). This reward mechanism successfully guided the agent to find circuits achieving chemical accuracy, validating RL’s feasibility. [129] introduced curriculum learning, enhancing the guidance mechanism towards chemical accuracy and demonstrating significant performance improvements over previous work. Following a similar RL-based paradigm, another line of research focused on applying these methods to variational quantum state diagonalization (VQSD). [135] proposed an RL-based approach to search for the optimal VQSD ansatz. While fundamentally similar to VQE, VQSD distinguishes itself by using the variance of the Hamiltonian as a loss instead of the energy expectation value. Built on this work, [136] introduced a mechanism based on quantum information to accelerate the guidance of the RL agent, achieving a speedup of at least two-fold. Further advancements were made in [137], where the authors replaced the multi-layer perceptron in RL with a Kolmogorov-Arnold Network, demonstrating enhanced performance in the search for effective ansatze.
III-C3 Sampling-based Learning Algorithms
@@
In addition to reinforcement learning, sampling-based learning algorithms represent another significant class of circuit synthesis methods. These algorithms initially define the search space, considering a quantum circuit with $n$ qubits, $m$ layers, and $k$ candidate gates, resulting in a search space of size $k^{m\times n}$. The methodology proceeds as follows: first, sampling is performed on the $n\times m$ grid points to generate a complete quantum circuit. Subsequently, the sampled circuit undergoes evaluation, informing modifications to the sampling strategy for future iterations. This process is repeated iteratively to complete circuit synthesis. Notably, sampling-based learning algorithms demonstrate compatibility with parameterized quantum gates, allowing for optimization of internal circuit parameters post-sampling. The primary challenge inherent in this approach stems from the discrete nature of the sampling process, which precludes the differentiable gradients. Consequently, developing effective mechanisms for updating the sampling strategy remains a critical area of research in this domain.
@@
[93] first employed Monte Carlo gradient estimation methods to update sampling strategies, demonstrating efficacy in designing Quantum Fourier Transform ansatz and mitigating circuit errors. [19] selected the optimal circuit through multiple sampling and ranking, testing their algorithms on four-qubit hydrogen ground state energy prediction and image classification QNN design problems. [110] proposed a method similar to [19], focusing on noise adaptation for quantum hardware. This approach generates shallow quantum circuits under the constraints of the quantum hardware to mitigate errors. Experimental results on IBM quantum devices demonstrated significant improvements over manually designed deeper circuits. However, both [19] and [110] utilized pre-defined circuit modules (termed super-circuits or supernets), substantially reducing the search space through pre-designed modules or specified layer arrangements. This approach, however, may limit the algorithms’ true capability, as their success heavily depends on these pre-designed circuit modules.
@@
[94] addressed this issue by questioning the efficacy of circuit structure sampling and synthesis and aligning algorithm applications. The study proposed using the arbitrary unitary approximation problem to benchmark QAS algorithms, considering this task the most challenging in QAS. To accommodate potentially deep decomposition depths, a simplified version - arbitrary circuit reconstruction - was introduced. This task provides a unitary matrix with a known circuit solution within a certain depth, enabling fairer algorithm comparisons. The results demonstrated that when prior information, e.g., pre-defined circuit modules, was removed, the performance of algorithms like [19] and [93] decreased significantly. This highlights their strong dependence on well-chosen modules. Notably, when candidate gates were limited to non-parameterized quantum gates, machine learning algorithms generally underperformed compared to heuristic algorithms.
@@
To enhance the efficiency and accuracy of searching algorithms, [20] proposed a method utilizing Gumbel-Softmax [138, 139, 140] to bridge the gap between discrete sampling and continuous gradient updates. This method uses the argmax term in the forward pass and the softmax term in backpropagation, addressing issues inherent in sampling-based learning algorithms and significantly improving search efficiency. The study eliminates prior knowledge interference by searching directly within the complete $k^{m\times n}$ space. It evaluates the method on three problems: VQE circuits (ground state energy prediction and Max-Cut) and QNN circuits (image classification), demonstrating notable improvements. Furthermore, the article examines the scalability of QAS algorithms, proposing two search processes: Macro Search and Micro Search. Macro Search conducts global domain searches, while Micro Search uses preset sub-circuit modules with shared structural parameters but distinct internal rotation parameters. Unlike [19] and [110], which predetermine circuit module structures, [20] searches for the module structures. They facilitate the identification of frequently reused modules, e.g., single and double excitation operators in ground state energy estimation or quantum convolution and pooling modules in image classification. This approach reduces the search space and enables the application of modules discovered in small-scale problems (e.g., six-qubit lithium hydride) to larger-scale problems (e.g., 18-qubit methane).
III-C4 Beyond Sampling-based
Given that the aforementioned papers [20, 93] render the entire search process differentiable end-to-end, we can also categorize these types of methods as differentiable quantum architecture search. Inspired by the rapid development of generative models, particularly denoising diffusion models [141, 142], one recent study [143] introduced a novel paradigm for quantum circuit synthesis. This approach bypasses the need for costly classical simulations of quantum dynamics by transforming the circuit synthesis problem into a generative one. Specifically, the authors encode quantum circuits into a continuous, high-dimensional tensor representation, allowing them to leverage powerful diffusion models to generate new circuits from noise. This method demonstrates an efficient and scalable way to explore the vast space of possible circuit configurations, offering a promising new direction for automated circuit design.  [119] proposed a neural network predictor for circuit synthesis, training it on small-scale quantum circuits and applying it to slightly larger-scale problems. This approach, while one of the few supervised learning models for QAS, faces limitations due to the diversity of quantum gates and the non-trivial relationship between gate configurations and circuit purposes. Recently, [130] introduced a training-free scheme based on [119], which reverted to sampling-based optimization. However, this study’s evaluation scale was limited (up to six qubits) and lacked comparison with state-of-the-art algorithms.
IV Quantum Logic Circuit Optimization
@@
Important factors need to be considered when executing quantum circuits, including qubit coherence time, environmental noise, etc. While quantum logic circuit synthesis generates circuits from scratch, optimization algorithms refine existing circuits based on specific targets. Consequently, circuits produced by synthesis algorithms can be further enhanced via optimization techniques. Moreover, by integrating the constraints of circuit synthesis with the optimization objectives, we can develop a unified synthesis optimization algorithm. It enables the direct creation of optimal quantum logic circuits that meet current optimization objectives.
IV-A Problem Definition
@@
We begin by presenting a formulation of the quantum logic circuit optimization problem. Analogous to the quantum logic circuit synthesis problem, we employ a unitary matrix to represent a given quantum circuit. Let $\mathbf{U}$ denote the unitary matrix of the original circuit to be optimized, and $\hat{\mathbf{U}}$ represent the unitary matrix of the optimized circuit. Given a set of optimization targets $\kappa_{i}$, where we assume that all targets are to be minimized, the corresponding objective is:
[…]
IV-B Quantum Logic Circuit Optimization Targets
@@
The objectives of quantum logic circuit optimization algorithms are directly correlated to their optimization methods. Table II illustrates the optimization methods in various research and their corresponding optimization objectives. It can be seen that, excluding brute-force search and reinforcement learning, other methods are aimed at more specific optimization targets. Overall, there is an emphasis on optimizing gate count and CNOT gates, while optimization of T gates is closely related to phase, resulting in a comparatively limited array of available optimization methods.
@@
TABLE II: Quantum logical circuit optimization methods versus optimization target.
@@ IV-B1 Gate Count
Optimization of the number of gates represents a fundamental objective in quantum circuit design. This optimization remains crucial across the current NISQ era and prospective FTQC paradigm. In the NISQ context, where qubit coherence time is limited and noise effects are significant, minimizing the number of quantum gates serves to reduce overall circuit noise, thereby enhancing operational accuracy. Moreover, in the realm of FTQC, gate count reduction contributes to improved computational efficiency of quantum hardware. Thus, the imperative to minimize quantum gates persists as a key strategy for advancing quantum circuit performance, irrespective of the underlying quantum computing framework. [144, 156, 28] guaranteed optimal Clifford circuits, with [28] demonstrating scalability. Works are targeting other gate sets, e.g., the parameterized gate set [145], Toffoli gates [54], Multi-Control Toffoli (MCT) circuits [53, 155], NCV circuits [55, 12], and parameterized VQE circuits [166].
IV-B2 Circuit Depth
@@
It is intrinsically associated with the total runtime of a quantum circuit, which is constrained by the coherence time of qubits. Consequently, optimizing circuit depth holds practical significance for scalable quantum computing in the NISQ era. While reducing gate count and minimizing circuit depth are often considered alike, many studies address both objectives concurrently [144, 29, 30]. However, these goals are not entirely identical. Depth optimization aims to enhance circuit parallelism, thereby reducing overall execution time and improving accuracy and efficiency.
@@
Circuit depth optimization is typically achieved by increasing the parallelism of quantum gates. Theoretically, a set of quantum gates without dependencies on each other can be executed in parallel. Dependencies in this context commonly refer to overlapping qubit operations. Thus, depth optimization can be understood as partitioning all gates in a circuit into the minimum number of mutually independent gate groups. Specifically,  [144] achieved optimal circuits through recursion, a method that can be used to construct optimal circuit datasets.  [167] proposed to transform the quantum circuit depth optimization problem into solving the edge coloring problem for lattice graphs. The algorithm presented in [62] extended beyond depth optimization with the capability of various goals. With the use of redundant qubits, [168, 169] have yielded bounds for the optimal circuit depth, serving as a critical benchmark for our optimization pursuits. [170, 171] further explored how these bounds change under different physical qubit connectivities, which is crucial for tailoring optimization strategies to specific hardware architectures. Other works focused on reducing the circuit depth of variational quantum circuits, e.g., QAOA circuits [172, 173, 174] and VQE circuits [175].
IV-B3 CNOT Count